What is the net present value of the following set of cash flows at a discount rate of 5 percent: Year 0 = -$62,000; Year 1 = $16,500; Year 2 = $23,800; Year 3 = $27,100 and Year 4 = $23,300?
Answer #1

The cashflows are:

C0 = -$62000

C1 = $16500

C2 = $23800

C3 = $27100

C4 = $23300

Discount rate = r = 5%

NPV is the present value of all the cash flows. NPV is calculated using the formula:

NPV = C0 + C1/(1+r)1 + C2/(1+r)2 + C3/(1+r)3 + C4/(1+r)4

NPV = -62000 + 16500/(1+5%)1 + 23800/(1+5%)2 + 27100/(1+5%)3 + 23300/(1+5%)4 = -62000 + 16500/1.051 + 23800/1.052 + 27100/1.053 + 23300/1.054 = -62000 + 15714.2857142857 + 21587.3015873016 + 23409.998920203 + 19168.9676626508 = 17880.5538844411 ~ 17880.55 (Rounded to two decimals)

Answer -> NPV = $17880.55